OpenStudy (anonymous):

What is the value for x in the proportion The quantity of x plus 8 all over the quantity of 5x minus 2 equals 3 over 8

OpenStudy (anonymous):

Not sure about the equation but still..... (x+8)/(5x-2)= 3/8 cross-multiply and solve for x :)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

idk how to find the answer

OpenStudy (anonymous):

U mean cross-multiplication part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its easy! U just have to multiply (5x-2) by 3 and (x+8) by 8. Give it a go and if u can't i'll do it for u :)

OpenStudy (anonymous):

ill try then reply

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

so 15x-6 and 8x+64 do I combine like terms?

OpenStudy (anonymous):

Yes and the equation would be something like this '15x-6= 8x-64'

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

Np :)